OpenStudy (anonymous):

I don't just want the answer... I want help on how to figure this out... I don't understand it pleasssseeee I will give medals The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) −1 −9 0 −1 1 7 g(x) g(x) = 3x − 2 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x).

OpenStudy (mathstudent55):

The slope of the line that passes through points \((x_1, y_1)\) and \((x_2, y_2)\) is \(m = \dfrac{y_2 - y_1}{x_2 - x_1} \)

OpenStudy (mathstudent55):

When the equation of a line is written in the slope-intercept form, \(y = mx + b\), m is the slope, and b is the y-coordinate of the y-intercept.

OpenStudy (anonymous):

right i have that so far -1 -(-9) over 0-(-1) =8

OpenStudy (mathstudent55):

To find the slope of function f(x), choose any to points from that function. I see that you are using the points (-1, -9) and (0, -1). \(m = \dfrac{-1 - (-9)}{0 - (-1)} = \dfrac{8}{1} = 8\) You are correct. the slope of function f(x) is 8. Now notice that function g(x) is written in the slope-intercept form, so you can read the slope right off the equation. What is the slope of the function g(x)?

OpenStudy (anonymous):

3

OpenStudy (mathstudent55):

Correct. The slope of f(x) is 8, and the slope of g(x) is 3. That takes care of part A.

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

I guess im confused they want a sentence how would i write that or am i reading into it to much :(

OpenStudy (mathstudent55):

I can tell you what to write, but it's already written above. Use your own calculation of the slope to show how you found the slope of f(x). Then state that the slope of g(x) is 3 because g(x) is written in the slope-intercept form, which allows you to read the slope right off the equation.

OpenStudy (anonymous):

oh ok sorry

OpenStudy (anonymous):

so for part b f(x) is steeper/Bigger than G(x) because the slope of F(x) is 8 and the slope of G(x) is only 3 ... is this correct

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Part A dealt with the slopes. Part B deals with the y-intercepts, not slopes.

OpenStudy (mathstudent55):

The y-intercept is the point where the line crosses the y-axis. It is b in the slope-intercept equation of a line.

OpenStudy (mathstudent55):

We know b for the g(x). It is -2 for g(x). Now we need to find b for f(x).

OpenStudy (mathstudent55):

Let's find the slope-intercept form of the equation of f(x) y = mx + b For f(x), we know m = 8 since we calculated it in part A. We have: y = 8x + b

OpenStudy (mathstudent55):

We need to find b. We can use any point in f(x) to find b. Let's use point (0, -1) y = 8x + b -1 = 8(0) + b -1 = b b = -1 That means the equation of f(x) is: f(x) = 8x - 1

OpenStudy (mathstudent55):

Now we see the y-intercept of f(x) is -1, and the y-intercept of g(x) is -2. Now we can compare them.

OpenStudy (anonymous):

ohh so F(x) is bigger because it is -1 and G9x) is -2

OpenStudy (anonymous):

that should say G(x) is -2

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

You are correct. You mean to say: The y-intercept of function f(x) is greater than the y-intercept of function g(x) because -1 is greater than -2.

OpenStudy (anonymous):

awesome thank you

OpenStudy (mathstudent55):

yw